# SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation

### v1

@@ 4.5 Cost Comparison of VLM Text Conversion
Table 7 compares the processing time of VLM-based text conversion between single-page processing and multiple semantic chunks processing. [^3] We used Qwen-2.5-VL-72B locally for this experiment. We can see that the average processing time for semantic chunks (our SCAN approach) is similar to the single-page processing time. The majority of input tokens are vision tokens, compared to the text tokens from conversion prompts. While the total input tokens for multiple chunks are fewer, the output tokens are more. This shows that our SCAN approach can reduce token-based costs (especially favorable for pay-as-token-usage API models) while extracting more textual information, all within approximately the same processing time.
@@ 5 Conclusion
We presented SCAN, a semantic document layout analysis approach for modern textual and visual RAG systems. By introducing coarse-grained semantic segmentation that preserves topical coherence, SCAN effectively reduces the information processing burden on VLMs while maintaining semantic integrity across document components. To develop SCAN, we labeled more than 24k document images with semantic layouts and trained a robust semantic layout analysis model. Our comprehensive evaluation across multiple datasets, languages, and document types demonstrates SCAN’s ability to enhance 1.7%–9.0% for textual RAG and 3.5%–6.4% for visual RAG. In addition, SCAN achieves these improvements while reducing token usage costs and maintaining the same processing time.
@@
In this work, we study semantic document layout analysis for RAG. To the best of our knowledge, there is no negative societal impact in this research. All our training data consists of publicly available PDFs from the internet, which similarly presents no ethical concerns. Our SCAN model aims to improve information extraction without introducing biases in the underlying content. We believe that improved document analysis can enhance accessibility of information for users across different languages and document formats. While our system improves RAG capabilities, users should still be mindful of the general limitations of AI systems when relying on generated answers.
We used Claude-3.7-Sonnet to polish the writing of the paper. We are responsible for all the materials presented in this work.
@@
1. Our SCAN model operates based on spatial image layout. In certain documents where contents that should logically form a single semantic chunk is physically separated in space and do not fit in a rectangular, our current model cannot yet establish these connections. This limitation could potentially be addressed through an additional trainable reading order model coupled with a semantic_box merging mechanism.
@@ Appendix A Examples from Our Semantic Layout Dataset
Figure 3 shows some examples of our semantic layout dataset. It contains diverse document pages, including research papers, administrative reports, user manuals, slides, flyers, and more. The dataset is annotated with the semantic_box class and the five global box classes: title, header, footer, date, and author.
Appendix B Training Details of Our Object Detection Models
@@
Appendix C Examples of Before and After the Predicted Box Selection
Figure 46 exemplify our predicted box selection strategy for the semantic document layout analysis.
Appendix D Experimental Details
D.1 Environment
For the text conversion with general VLMs, we used 8 NVIDIA L40S (48GB) GPUs with INTEL(R) XEON(R) GOLD 6548N CPU, and the details are as follows.
@@ D.1 Environment
D.2 Prompts for the VLM Text Conversion
@@ 2. Formula Processing:
[…]
@@ D.3 Prompts for LLM-as-a-judge
Figure 3: Examples of our semantic layout dataset.
Hyper-parameter YOLO11-X RT-DETR-X
Batch size {8, 16, 32, 64} {8, 16, 32, 64}
Learning rate {1e-4, 5e-4, 1e-3, 5e-3} {5e-5, 1e-4, 5e-4, 1e-3, 5e-3}
Max training epochs {30, 40, 80, 120} {80, 120, 160}
Weight decay {5e-5, 1e-4, 5e-4} {1e-5, 1e-4, 1e-3}
Warmup epochs {5, 10} {5, 10}
Image size 1024
Dropout 0.0
Optimizer AdamW
Learning rate scheduler cos_lr
Figure 4: Semantic document layout analysis result before (left) and after (right) our predicted box selection strategy (1/3).
Figure 5: Semantic document layout analysis result before (left) and after (right) our predicted box selection strategy (2/3).
Figure 6: Semantic document layout analysis result before (left) and after (right) our predicted box selection strategy (3/3).

### v2

@@ 4.4 Cost Comparison of VLM Text Conversion
Our approach, which splits a page into multiple images, consistently improves the accuracy of textual RAG. However, one concern is that increasing the number of images to be processed may also increase processing time and cost. To investigate this, we compared the processing time and token usage of VLM-based text conversion with and without applying SCAN.
Chunking Method Model Architecture (# Parameters) # Chunks Per Image Relative Chunk Size (%) Textual RAG Score Visual RAG Score
No chunking N/A 1.0 100.0 31.1 70.2
DiT DiT (304M) 12.3 16.3 33.5 64.9
DocLayout-YOLO YOLO11-X (56.9M) 9.9 11.3 22.4 59.6
Beehive RT-DETR-L (42.8M) 17.4 4.8 25.3 65.8
[…]
We randomly sampled 10 images from OHR-Bench and used Qwen-2.5-VL-72B for text conversion. [^3] We used a vLLM vllm server running on four NVIDIA RTX PRO 6000 Blackwell GPUs for efficient batch processing. Table 10 reports the comparison between single-page processing and multiple semantic chunk processing. The results show that both the number of input tokens and output tokens increase when using multiple semantic chunks. This implies that applying SCAN with API-based models that charge by token usage may lead to higher monetary cost. However, despite the increase in token counts, the average processing time is reduced. We attribute this to the substantial decrease in the number of input tokens per request to the VLM, which lowers the cost of attention computation. Specifically, while the average number of input tokens per chunk is 1,320.4 in the single-page setting, it is reduced to 780.9 (=9683.1/12.4=9683.1/12.4) in the multiple-chunk setting. Thus, SCAN not only enables the extraction of richer textual information but also reduces computational overhead.
4.5 Chunking Granularity and RAG Performance
To quantitatively evaluate how chunking granularity affects RAG performance, we computed the average number of chunks per image and their relative sizes for the five different chunking methods (Table 11). From the number and relative size of chunks, we observe that our SCAN divides a page into relatively larger and fewer chunks compared to the other chunking methods. SCAN achieves the highest performance in both textual and visual RAG, indicating that its moderate level of granularity is well-suited for RAG tasks.
[…]
@@ 5 Conclusion
We presented SCAN, a semantic document layout analysis approach for modern textual and visual RAG systems. By introducing coarse-grained semantic segmentation that preserves topical coherence, SCAN effectively reduces the information processing burden on VLMs while maintaining semantic integrity across document components. To develop SCAN, we labeled more than 24k document images with semantic layouts and trained a robust semantic layout analysis model. Our comprehensive evaluation across multiple datasets, languages, and document types demonstrates SCAN’s ability to enhance textual and visual RAG performance by 2.7–9.4 and 5.6–10.4 points, respectively. In addition, SCAN achieves these improvements while reducing computational costs.
Future work could explore SCAN’s applicability to other document understanding tasks, such as document summarization, information extraction, and document VQA.
@@
In this work, we study semantic document layout analysis for RAG. To the best of our knowledge, there is no negative societal impact in this research. All our training data consist of publicly available PDFs from the internet, which similarly present no ethical concerns. Our SCAN model aims to improve information extraction without introducing biases in the underlying content. We believe that improved document analysis can enhance the accessibility of information for users across different languages and document formats. While our system improves RAG capabilities, users should still be mindful of the general limitations of AI systems when relying on generated answers.
We used Claude-3.7-Sonnet and GPT-5 to help polish the writing of the paper. We are responsible for all the materials presented in this work.
@@
1. Our SCAN model operates based on spatial image layout. In certain documents where contents that should logically form a single semantic chunk is physically separated in space and do not fit in a rectangular box, our current model cannot yet establish these connections. This limitation could potentially be addressed through an additional trainable reading order model coupled with a semantic box merging mechanism.
@@ Appendix A Examples from Our Semantic Layout Dataset
Figure 3 shows some examples of our semantic layout dataset. It contains diverse document pages, including research papers, administrative reports, user manuals, slides, flyers, and more. The dataset is annotated with the semantic box class and the five global box classes: title, header, footer, date, and author.
Figure 3: Examples of our semantic layout dataset.
Appendix B Annotation Instructions for Our Semantic Layout Dataset
This section summarizes the guideline we provided to annotators when creating our semantic layout dataset. Annotators first locate and annotate page-level global boxes such as headers and footers. After that, they segment the remaining content into subtopics so that each subtopic is as semantically self-contained as possible. For example, in a product catalog (e.g., the bottom-left example in Figure 3), the description of each product is independent of the others and therefore receives its own semantic box. If a single subtopic still contains too much information, annotators further split it so that each semantic box contains a manageable amount of content, which is suitable for processing by VLMs.
Semantic units are always annotated with axis-aligned rectangles. Rectangles are required not to overlap and, taken together, should cover all content in the page. When a single rectangle would inevitably contain content from multiple topics, the region is split so that each rectangle corresponds to a single topic. For instance, in the top-right example in Figure 3, the bottom-right semantic box is topically related to the upper box, but they are annotated separately because a single rectangle would also include the bottom-left region, which belongs to a different topic.
During annotation, we discard pages that are unsuitable for our task. Typical exclusion cases include:
  • pages written in languages other than Japanese or with text so small that the content cannot be reliably read
  • pages that are heavily rotated or not displayed in the correct orientation
Appendix C Training Details of Our Object Detection Models
@@
Hyper-parameter YOLO11-X RT-DETR-X
Batch size {8, 16, 32, 64} {8, 16, 32, 64}
Learning rate {1e-4, 5e-4, 1e-3, 5e-3} {5e-5, 1e-4, 5e-4, 1e-3, 5e-3}
Max training epochs {30, 40, 80, 120} {80, 120, 160}
Weight decay {5e-5, 1e-4, 5e-4} {1e-5, 1e-4, 1e-3}
Warmup epochs {5, 10} {5, 10}
Image size 1024
Dropout 0.0
Optimizer AdamW
Learning rate scheduler cos_lr
Appendix D Output Examples of Our Semantic Layout Analysis Model
This section presents qualitative examples of SCAN predictions. Figures 48 illustrate that SCAN can handle complex layouts in both Japanese and English documents, grouping related elements into coherent semantic chunks.
Figures 7 and 8 also reveal typical failure patterns, where some regions are covered by overlapping boxes. Such redundancy is usually harmless when semantic chunks are used as retrieval units in RAG or as inputs to LLMs, but it may be undesirable in traditional OCR pipelines that aim to extract each character exactly once. For these use cases, it would be beneficial to add post-processing that merges or removes overlapping boxes.
Figure 4: Examples of SCAN model outputs.
Figure 5: Example of SCAN model outputs.
Figure 6: Example of SCAN model outputs.
Figure 7: Example of SCAN model outputs.
Figure 8: Example of SCAN model outputs.
Appendix E Experimental Details
E.1 Environment
For text conversion with general VLMs, we used 8 NVIDIA L40S (48GB) GPUs with INTEL(R) XEON(R) GOLD 6548N CPU, and the details are as follows.
@@ E.1 Environment
E.2 Prompts for the VLM Text Conversion
@@ 2. Formula Processing:
- Enclose block formulas with \$\$ \$\$. For example: \$\$ $\frac${-b $\pm$ $\sqrt${b^2 - 4ac}}{2a} \$\$.